# Double rotatory power reversal and continuous Kerr angle in bi-isotropic media with anomalous Hall current

### v1

@@ Abstract
with $\Delta_{\pm,\pm}^{\prime}=\operatorname{Re}[\Delta_{\pm,\pm}]$ and $\Delta_{\pm,\pm}^{\prime\prime}=\operatorname{Im}[\Delta_{\pm,\pm}]$. The real and imaginary parts provide the Kerr angles (ellipticity and rotation) [31, 8], respectively,
@@
[…]
Recalling the refractive indices in Eq. (4), we note that $n_{LCP,+}$ and $n_{RCP,+}$ are connected to the usual refraction. On the other hand, the indices $n_{LCP,-}$ and $n_{RCP,-}$ (for $\omega>\omega_{R}$) are related to negative refraction. Thus, the element $\Delta_{+,+}$ provides the Kerr angles for the usual refraction case, $\Delta_{-,-}$ yields the Kerr angles for the pure negative refraction scenario, and $\Delta_{\pm,\mp}$ are associated with mixing effects of normal and negative refraction.
Figure 4: Kerr rotation $\theta_{K}^{+,\pm}$ of Eq. (15). The red line shows the behavior of $\theta_{K}^{+,+}$, composed by the indices $n_{LCP,+}$ and $n_{RCP,+}$. The blue line depicts $\theta_{K}^{+,-}$, constituted with $n_{LCP,+}$ and $n_{RCP,-}$. The purple line is defined for the region where both $\theta_{K}^{+,\pm}$ are null. Here, we have used $n_{1}=1$, $\mu_{1}=1$ $\mu=1$, $\epsilon=3,\alpha^{\prime\prime}=3$, and $b=1$ $\mathrm{s}^{-1}$.
As previously discussed, absorption occurs only in the range $\omega<\omega_{0}$, being associated with RCP refractive indices. Thus, only in this frequency interval, one finds $\Delta^{\prime\prime}\neq 0$, providing a non-null Kerr rotation for the reflected wave. In the following subsection, we examine the Kerr rotation angles for $\Delta_{+,\pm}$ and $\Delta_{-,\pm}$ in terms of frequency.
The general behaviors of $\theta_{K}^{+,\pm}$ in terms of the frequency are illustrated in Fig. 4. We observe that the factors $\theta_{Kerr}^{+,+}$ and $\theta^{+,-}_{K}$ are positive and negative, respectively, meaning that the polarization ellipse of the reflected wave is rotated in the counterclockwise (clockwise) direction relative to $x$-axis of polarization basis [45].
The continuous Kerr rotation angle of Fig. 4 is a remarkable peculiar characteristic of this bi-isotropic system endowed with AHE term, in contrast to the usual discontinuous Kerr rotation angles that occur in media ruled by the axion electrodynamics [11]. For the dielectric system here considered, such a discontinuity takes place only when
$n_{1}>\mu_{1}\alpha^{\prime\prime},$ (17)
for which the squared magnitude of coefficient $\Delta_{+,+}$ becomes equal to unity, leading to a divergence for the Kerr rotation angle tangent ($\theta_{K}=\pi/4$). The latter manifests itself as an abrupt sign change (discontinuity) in $\theta_{Kerr}^{+,+}$ at the frequency $\tilde{\omega}$ given by $\tilde{\omega}={\mu_{1}^{2}b}/({\mu_{1}^{2}\epsilon+\mu n_{1}^{2}})$. Figure 5 illustrates the frequency behavior of $\theta_{Kerr}^{+,+}$ under condition (17). Notice that at the frequencies $\tilde{\omega}_{1}$ and $\tilde{\omega}_{2}$ for the red (blue) line, respectively, the rotation of the polarization ellipse changes abruptly, indicating a reversal in the ellipse polarization rotation.
Figure 5: Kerr rotation $\theta_{K}^{+,+}$ of Eq. (15) under the condition (17). The red and blue lines show $\theta_{K}^{+,+}$ (obtained by using $n_{LCP,+}$ and $n_{RCP,+}$) for $n_{1}=1.5$ and $n_{1}=1.9$, respectively. The purple line indicates $\theta_{K}^{+,+}=0$. Here, we have used $\mu_{1}=1$ $\mu=1$, $\epsilon=3,\alpha^{\prime\prime}=1$, $b=1$ $\mathrm{s}^{-1}$, $n_{1}=1.5$ (red), and $n_{1}=1.9$ (blue).
The other Kerr rotation angles, $\theta_{K}^{-,\pm}$, which are obtained by using $n_{LCP,-}$, can be easily found by implementing $\Delta_{-,\pm}=\left.\Delta_{+,\pm}\right|_{Q_{+}\rightarrow-Q_{+}}$, and will not be addressed here.
Figure 6: Kerr ellipticity $\eta_{K}^{+,\pm}$ of Eq. (16). The red line shows $\eta^{+,+}_{K}$, while the blue curve depicts $\eta^{+,-}_{K}$. The purple line indicates the region both $\eta_{K}^{+,\pm}$ are equal to each other. Here, we have used $n_{1}=1$, $\mu_{1}=1$, $\mu=1,\epsilon=3,\alpha^{\prime\prime}=3$, and $b=1$ $\mathrm{s}^{-1}$. When $\eta_{K}=0$, the reflected wave is linearly polarized. The extreme values (±$\pm\pi/4$) are associated with left- and right-handed circularly polarized waves.
Kerr ellipticity angles $\eta_{K}^{\pm,\pm}$. The Kerr ellipticities, given by Eq. (16), are non-null for any frequency since the refractive indices (4) possess at least one real piece. For $0<\omega<\omega_{0}$, it holds $\operatorname{Re}[n_{RCP,\pm}]=\mu\alpha^{\prime\prime}$, so that the four initial possibilities of Eq. (16) are reduced to only two, in accordance with the double sign of $n_{LCP,\pm}$. For the index $n_{LCP,+}$, we define $\eta_{K}^{+,\pm}$, whose behavior in terms of frequency is illustrated in Fig. 6. For $\omega<\omega_{0}$, one has $\eta^{+,+}_{K}=\eta^{+,-}_{K}$, as already explained, behavior represented by the purple line in Fig. 6. In the frequencies $\omega_{\pm}$, given by
the ellipticity is ±$\pm\pi/4$, respectively, where the wave is circularly polarized. In the range $\omega>\omega_{0}$, the indices $n_{RCP,\pm}$ are real, assuming two different values, therefore making $\eta^{+,+}_{K}\neq\eta^{+,-}_{K}$, splitting the Kerr ellipticity into two distinct lines (red and blue) in Fig. 6.
Reflectance and super reflectance effect. Recently, anomalous reflection ($R>1$), at normal incidence, has been reported as one key feature of Weyl semimetals in the presence of both axion terms ($b_{0}$ and $\mathbf{b}$) [38]. Here, we report that a similar effect can appear in bi-isotropic media modified by the AHE term. Starting from the Fresnel coefficients (12a) and (12b), one obtains the reflection coefficients,
$R_{LCP}^{\pm}$ $=\left|\dfrac{\mu_{1}\left(\mu\alpha^{\prime\prime}\mp Q_{+}\right)+\mu n_{1}}{\mu_{1}\left(\mu\alpha^{\prime\prime}\mp Q_{+}\right)-\mu n_{1}}\right|^{2},$ (19a)
$R_{RCP}^{\pm}$ $=\left|\dfrac{\mu_{1}\left(\mu\alpha^{\prime\prime}\pm Q_{-}\right)-\mu n_{1}}{\mu_{1}\left(\mu\alpha^{\prime\prime}\pm Q_{-}\right)+\mu n_{1}}\right|^{2},$ (19b)
where $Q_{\pm}$ are given in Eqs. (5). We then address the two reflection amplitudes, $R_{LCP}^{-}$ and $R_{RCP}^{-}$, stemming from $n_{LCP,-}$ (always negative for all frequencies) and $n_{RCP,-}$ (with negative real part for $\omega>\omega_{R}$). In Fig. 7, we illustrate reflectances that exhibit anomalous behaviors. As for $R_{RCP}^{-}$ (red line), it initially decreases with frequency, reaches a null value near $\omega_{0}$, and then augments monotonically, becoming greater than unity for $\omega>\omega_{R}$, which is exactly the frequency whereupon the refractive index $n_{RCP,-}$ becomes negative.
Figure 7: Reflection coefficients $R_{RCP}^{-}$ (red line) and $R_{LCP}^{-}$ (blue line). $R_{RCP}^{-}$ becomes greater than unit for $\omega>\omega_{R}$. $R_{LCP,-}$ (blue line) exhibits ubiquitous superreflectance. The inset plot highlights the behavior near the frequency $\omega_{0}$, below which $n_{RCP,-}$ has a non-null imaginary part. Here, we have used $\mu_{1}=1,n_{1}=\sqrt{2}$, $\mu=1,\epsilon=3,\alpha^{\prime\prime}=3$, $b=1$ $\mathrm{s}^{-1}$, which implies $\omega_{0}=0.1$ and $\omega_{R}=1/3$ $\mathrm{s}^{-1}$.
On the other hand, the coefficient $R_{LCP}^{-}$ (blue line) already starts at unit, $R_{LCP}^{-}=1$, and progressively enlarges with the frequency, defining an increasing anomalous reflectance ($R$ greater than unit) for the entire frequency domain. Both $R_{LCP}^{-}$ and $R_{RCP}^{-}$ are endowed with super reflectance, being this property associated with negative refraction, a hallmark of metamaterials [44, 5]. The negative refraction of the indices $n_{RCP}^{-}$ and $n_{LCP}^{-}$ occurs as a consequence of an anomalous increasing wave magnitude that propagates along the direction $-\hat{z}$ [38].
Final remarks. In this work, we investigated the optical properties of bi-isotropic materials endowed with the anomalous Hall current of the axion electrodynamics. Four dispersive and distinct refractive indices were obtained, implying richer optical effects, such as rotatory power, Kerr rotation/ellipticity, and reflectance. Birefringence aspects were examined in terms of four rotatory power coefficients. An RP endowed with double sign reversal was reported, being a remarkable characteristic of this particular dispersive bi-isotropic media with the AHE. The first signal reversal occurs for $\omega<\omega_{0}$, while the second one happens for $\omega>\omega_{0}$, the range in which the contribution of the full $n_{\text{RCP},-}$ index emerges, as shown in Fig. 3.
We also considered an interface separating two media: dielectric medium 1 ($\epsilon_{1}$,$\mu_{1}$) and bi-isotropic medium 2 with AHE ($\epsilon$,$\mu$, $b$). For normal incidence, we have written four distinct Kerr rotation angles, see Eq. (15). A continuous Kerr rotation angle (deprived of discontinuity) was found, as shown in Fig. 4, providing another key optical feature of bi-isotropic media with AHE. Such behavior occurs whenever the inequality (17) does not hold. In typical Weyl semimetals scenarios, the absence of the magnetoelectric parameter, $\alpha^{\prime\prime}=0$, assures the condition (17) as granted, in such a way the Kerr angle discontinuity is observed. The Kerr ellipticity was also analyzed, with its dispersive behavior illustrated in Fig. 6. The reflected wave can exhibit linear polarization ($\eta=0$) at two distinct frequencies, left- and right-handed circular polarization ($\eta=\pm\pi/4$), and elliptical polarization (other values). Finally, the reflectance for normal incidence was also considered, starting from the Fresnel coefficients in Eq. (12). Four reflectances were carried out, with two of them depicted in Fig. 7, where an anomalous reflection behavior ($R>1$) becomes evident at all frequencies for the blue line ($n_{\text{LCP},-}$) and for $\omega>\omega_{R}$ ($n_{\text{RCP},-}$), the exact regions where these refractive indices are negative. This effect in chiral media was recently reported in the literature for Weyl semimetals [38] when some electromagnetic parameters are negative. For the enriched bi-isotropic medium here examined, however, the anomalous reflectance can occur taking as positive all the involved parameters.

### v2

@@ IV Kerr rotation and Kerr ellipticity
with $\Delta_{+,\pm}^{\prime}=\operatorname{Re}[\Delta_{+,\pm}]$ and $\Delta_{+,\pm}^{\prime\prime}=\operatorname{Im}[\Delta_{+,\pm}]$. For $\Delta_{-,-}$, one writes analogous expressions. The real and imaginary parts provide the Kerr angles (ellipticity and rotation) [42, 14], respectively,
$\tan(\,2\theta_{K}^{+,\pm}\,)$ $=-\frac{2\Delta_{+,\pm}^{\prime\prime}}{1-\left|\Delta_{+,\pm}\right|^{2}},$ (16)
$\sin(\,2\eta_{K}^{+,\pm}\,)$ $=\frac{2\Delta_{+,\pm}^{\prime}}{1+\left|\Delta_{+,\pm}\right|^{2}}.$ (17)
Recalling the refractive indices in Eq. (4), we note that $n_{2,+}$ and $n_{1,+}$ are connected to the usual refraction (for $\omega<\omega_{R}$). On the other hand, the indices $n_{2,-}$ and $n_{1,-}$ (for $\omega>\omega_{R}$) are related to negative refraction. Thus, the element $\Delta_{+,+}$ provides the Kerr angles for the usual refraction case, $\Delta_{-,-}$ yields the Kerr angles for the pure negative refraction scenario for ($\omega>\omega_{R}$).
Figure 4: Kerr rotation $\theta_{K}^{+,\pm}$ of Eq. (16). The red line shows the behavior of $\theta_{K}^{+,+}$, composed by the indices $n_{2,+}$ and $n_{1,+}$. The blue line depicts $\theta_{K}^{+,-}$, constituted with $n_{2,+}$ and $n_{1,-}$. The purple line is defined for the region where both $\theta_{K}^{+,\pm}$ are null. Here, we have used $n_{1}=1$, $\mu_{1}=1$ $\mu=1$, $\epsilon=3,\alpha^{\prime\prime}=3$, and $b=1$ $\mathrm{s}^{-1}$.
As previously discussed, absorption occurs only in the range $\omega<\omega_{0}$, being associated with $n_{1,\pm}$ refractive indices. Thus, only in this frequency interval, one finds $\Delta^{\prime\prime}\neq 0$, providing a non-null Kerr rotation for the reflected wave. In the following subsection, we examine the Kerr rotation angles for $\Delta_{+,\pm}$ in terms of frequency.
The general behaviors of $\theta_{K}^{+,\pm}$ in terms of the frequency are illustrated in Fig. 4. We observe that the factors $\theta_{K}^{+,+}$ and $\theta^{+,-}_{K}$ are positive and negative, respectively, meaning that the polarization ellipse of the reflected wave is rotated in the counterclockwise (clockwise) direction relative to $x$-axis of polarization basis [56].
The giant continuous Kerr rotation angle of Fig. 4 can be close to the magnitude $\pi/18$ and represents a remarkable peculiar characteristic of this bi-isotropic system endowed with AHE term, in contrast to the usual discontinuous Kerr rotation angles that occur in media ruled by the axion electrodynamics [17]. For the dielectric system considered here, such a discontinuity takes place only when
$n_{1}>\mu_{1}\alpha^{\prime\prime},$ (18)
for which the squared magnitude of coefficient $\Delta_{+,+}$ becomes equal to unity, leading to a divergence for the Kerr rotation angle tangent ($\theta_{K}=\pi/4$). The latter manifests itself as an abrupt sign change (discontinuity) in $\theta_{K}^{+,+}$ at the frequency $\tilde{\omega}$ given by $\tilde{\omega}={\mu_{1}^{2}b}/({\mu_{1}^{2}\epsilon+\mu n_{1}^{2}})$.
Figure 5 illustrates the frequency behavior of $\theta_{Kerr}^{+,+}$ under condition (18). Notice that at the frequencies $\tilde{\omega}_{1}$ and $\tilde{\omega}_{2}$ for the red (blue) line, respectively, the rotation of the polarization ellipse changes abruptly, indicating a reversal in the ellipse polarization rotation. The other Kerr rotation angles, $\theta_{K}^{-,-}$, which are obtained by using $n_{2,-}$, can be easily found by implementing $\Delta_{-,-}=\left.\Delta_{+,+}\right|_{Q_{+}\rightarrow-Q_{+}}$, and will not be addressed here.
Figure 5: Kerr rotation $\theta_{K}^{+,+}$ of Eq. (16) under the condition (18). The red and blue lines show $\theta_{K}^{+,+}$ (obtained by using $n_{2,+}$ and $n_{1,+}$) for $n_{1}=1.5$ and $n_{1}=1.9$, respectively. The purple line indicates $\theta_{K}^{+,+}=0$. Here, we have used $\mu_{1}=1$ $\mu=1$, $\epsilon=3,\alpha^{\prime\prime}=1$, $b=1$ $\mathrm{s}^{-1}$, $n_{1}=1.5$ (red), and $n_{1}=1.9$ (blue).
Figure 6: Kerr ellipticity $\eta_{K}^{+,\pm}$ of Eq. (17). The red line shows $\eta^{+,+}_{K}$, while the blue curve depicts $\eta^{+,-}_{K}$. The purple line indicates the region both $\eta_{K}^{+,\pm}$ are equal to each other. Here, we have used $n_{1}=1$, $\mu_{1}=1$, $\mu=1,\epsilon=3,\alpha^{\prime\prime}=3$, and $b=1$ $\mathrm{s}^{-1}$. When $\eta_{K}=0$, the reflected wave is linearly polarized. The extreme values ($\pm\pi/4$) are associated with left- and right-handed circularly polarized waves.
V Kerr ellipticity angles $\eta_{K}^{+,\pm}$
The Kerr ellipticities, given by Eq. (17), are non-null for any frequency since the refractive indices (4) possess at least one real piece. For $0<\omega<\omega_{0}$, it holds $\operatorname{Re}[n_{1,\pm}]=\mu\alpha^{\prime\prime}$, so that the initial possibilities of Eq. (17) are reduced to only one. For the index $n_{2,+}$, we define $\eta_{K}^{+,\pm}$, whose behavior in terms of frequency is illustrated in Fig. 6. For $\omega<\omega_{0}$, one has $\eta^{+,+}_{K}=\eta^{+,-}_{K}$, as already explained, behavior represented by the purple line in Fig. 6. In the range $\omega>\omega_{0}$, the indices $n_{1,\pm}$ are real, assuming two different values, therefore making $\eta^{+,+}_{K}\neq\eta^{+,-}_{K}$, splitting the Kerr ellipticity into two distinct lines (red and blue, Fig. 6). In the frequencies $\omega_{\pm}$, given by
the ellipticity is $\pm\pi/4$, respectively, where the wave is circularly polarized. As mentioned previously, $\Delta_{+,-}$ is only defined for $\omega<\omega_{R}$, being thus important to ensure that the frequencies $\omega_{+}$ and $\omega_{-}$ are within this range [^2] The frequencies $\omega_{+}$ and $\omega_{-}$ are within the range $\omega<\omega_{R}$ when (i) $2\alpha^{\prime\prime}\sqrt{\mu_{1}\epsilon_{1}}>\epsilon_{1}$, for $\omega_{+}<\omega_{R}$; and (ii) $2\alpha^{\prime\prime}\sqrt{\mu_{1}\epsilon_{1}}>2\frac{\mu_{1}}{\mu}\epsilon-\epsilon_{1}$, for $\omega_{-}<\omega_{R}$; being it the case of Fig. 6., as shown in Fig. 6.
VI Reflectance and super reflectance effect
Recently, anomalous reflection ($R>1$), at normal incidence, has been reported as one key feature of Weyl semimetals in the presence of both axion terms ($b_{0}$ and $\mathbf{b}$) [49]. Here, we report that a similar effect can appear in bi-isotropic media modified by the AHE term. Starting from the Fresnel coefficients (12a) and (12b), one obtains the reflection coefficients,
$R_{2}^{\pm}$ $=\left|\dfrac{\mu_{1}\left(\mu\alpha^{\prime\prime}\mp Q_{+}\right)+\mu n_{1}}{\mu_{1}\left(\mu\alpha^{\prime\prime}\mp Q_{+}\right)-\mu n_{1}}\right|^{2},$ (20a)
$R_{1}^{\pm}$ $=\left|\dfrac{\mu_{1}\left(\mu\alpha^{\prime\prime}\pm Q_{-}\right)-\mu n_{1}}{\mu_{1}\left(\mu\alpha^{\prime\prime}\pm Q_{-}\right)+\mu n_{1}}\right|^{2},$ (20b)
where $Q_{\pm}$ are given in Eqs. (5). We then address the two reflection amplitudes, $R_{1}^{+}$ and $R_{1}^{-}$, stemming from $n_{1,+}$ (always positive for all frequencies) and $n_{1,-}$ (with negative real part for $\omega>\omega_{R}$). In Fig. 7, we illustrate reflectances that exhibit anomalous behaviors. The coefficient $R_{1}^{+}$ (blue line) already starts at unit, $R_{1}^{+}=1$, progressively diminishes with the frequency, reaching a mininimum near the frequency $\omega_{0}$, and then monotonically increases with frequency to its assymptotical value. On the other hand, $R_{1}^{-}$ (red line) initially decreases with frequency, reaches a null value near $\omega_{0}$, and then augments monotonically, becoming greater than unity for $\omega>\omega_{R}$, which is exactly the frequency whereupon the refractive index $n_{1,-}$ becomes negative. Thus, $R_{1}^{-}$ is endowed with super reflectance, being this property associated with negative refraction, a hallmark of metamaterials [57, 5]. The negative refraction of the index $n_{1}^{-}$ occurs as a consequence of an anomalous increasing wave magnitude that propagates along the direction $\hat{z}$ [49].
Figure 7: Reflection coefficients $R_{1}^{-}$ (red line) and $R_{1}^{+}$ (blue line). $R_{1}^{-}$ becomes greater than unit for $\omega>\omega_{R}$. The inset plot highlights the behavior near the frequency $\omega_{0}$, below which both $n_{1,\pm}$ have non-null imaginary parts. The solid purple line indicates that $R_{1}^{+}$ and $R_{1}^{-}$ are approximately equal to each other near the origin. Here, we have used $\mu_{1}=1,n_{1}=\sqrt{2}$, $\mu=1,\epsilon=3,\alpha^{\prime\prime}=3$, $b=1$ $\mathrm{s}^{-1}$, which implies $\omega_{0}=0.1$ and $\omega_{R}=1/3$ $\mathrm{s}^{-1}$.
VII Final remarks
In this work, we investigated the optical properties of bi-isotropic materials endowed with the anomalous Hall current of axion electrodynamics. Four dispersive and distinct refractive indices were obtained, implying richer optical effects, such as rotatory power, Kerr rotation/ellipticity, and reflectance. Birefringence aspects were examined in terms of three distinct rotatory power coefficients. An RP endowed with double sign reversal was reported, being a remarkable characteristic of this particular dispersive bi-isotropic media with the AHE. The first signal reversal occurs for $\omega<\omega_{0}$, while the second one happens for $\omega>\omega_{0}$, the range in which the contribution of the full $n_{\text{1},-}$ index emerges, as shown in Fig. 3.
We also considered an interface separating two media: dielectric medium 1 ($\epsilon_{1}$$\mu_{1}$) and bi-isotropic medium 2 with AHE ($\epsilon$, $\mu$, $b$). For normal incidence, we have written three distinct Kerr rotation angles; see Eq. (16). A continuous Kerr rotation angle (deprived of discontinuity) was found, reaching giant magnitudes as large as $\pi/18$, see Fig. 4, providing another key optical feature of bi-isotropic media with AHE. Such behavior occurs whenever the inequality (18) does not hold. In typical Weyl semimetals scenarios, the absence of the magnetoelectric parameter, $\alpha^{\prime\prime}=0$, assures the condition (18) as granted, in such a way that the Kerr angle discontinuity is observed. The Kerr ellipticity was also analyzed, with its dispersive behavior illustrated in Fig. 6. The reflected wave can exhibit linear polarization ($\eta=0$) at two distinct frequencies, left- and right-handed circular polarization ($\eta=\pm\pi/4$), and elliptical polarization (other values). Finally, the reflectance for normal incidence was also considered, starting from the Fresnel coefficients in Eq. (12). Four reflectances were carried out, with two of them depicted in Fig. 7, where an anomalous reflection behavior ($R>1$) becomes evident in the range $\omega>\omega_{R}$ for the red line ($n_{1,-}$), the exact region where this refractive index becomes negative. Super reflectance ($R>1$) was observed in a special class of materials (amplifying matter) under conditions of enhanced total reflection [58], in which the refractive index of medium 1 is greater than that of medium 2 [59]. In our case, the coefficient $R>1$ happens as a consequence of the negative refractive index that occurs in certain frequency ranges, for which the refractive index of medium 2 becomes smaller than that of medium 1. In this sense, the phenomenon we have reported is analogous (at least in an effective way) to the enhanced reflection described in the 1970s. Furthermore, the increasing reflectance shown in Fig. 7 is dispersive, differing from the non-dispersive scenario of Refs. [58, 59]. A previously suggested explanation for $R>1$ attributes this effect to mechanisms occurring in medium 2 that supply energy to the reflected wave  [59, 60]. In Weyl semimetals, with both chiral magnetic conductivity and AHE, the chiral imbalance drives the system out of equilibrium, and relaxation towards the equilibrium can supply the additional energy involved in the anomalous reflectance [48, 49].
@@ VII Final remarks
Materials with bi-isotropic constitutive relations (or isotropic magnetoelectric tensors) can be found in the class of pyrochlores [62], such as Eu2Ir2O7 and Nd2Ir2O7. For both, anomalous Hall conductivity has been reported. See Ref. [63] and [64]. Other notable bi-isotropic materials include the chalcogenides Bi2Se3, Bi2Te3, and Sb2Te3 [65]. In particular, AHE has been observed in Bi2Se3, as reported in Refs. [66, 67]. In principle, bi-isotropic compounds in the presence of the AHE and non-equilibrium configurations gather the conditions for manifesting the effects discussed in this work, opening up exciting new avenues for further investigations on chiral media, as already done for Weyl semimetals in Refs. [48, 49, 50].
Acknowledgments